### v3

@@ 3 A Dynamic Regional Model of Climate and the Economy with an ETS
We now introduce a dynamic regional model of climate and the economy that integrates a global ETS between 12 aggregated regions. Our model framework extends the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps. The main objective of this study is to present a potential dynamic path of carbon emission prices in a noncooperative environment where each region aims to maximize its own social welfare by choosing the optimal amount of emission abatement, emission permits purchased from the global permits market, and the amount of its consumption. The model is based on the assumption that future regional emissions are strictly constrained by the emission caps set by the commitments under the Paris Agreement (and revised under the Glasgow Pact) and net zero emissions targets. In this study, we consider the case without uncertainty and without transaction cost in the global ETS.
The macroeconomic framework of our model employs a multi-regional representation of the Ramsey growth model. We follow the RICE model and assume that the world consists of 12 aggregated regions: the United States (US), the EU, Japan, Russia, Eurasia, China, India, Middle East (MidEast), Africa, Latin America (LatAm), Other High-Income countries (OHI) and other non-OECD Asia (OthAs). [^8] See Appendix 1 A.2 for the full list of countries and regional aggregation. Each region is indexed by $i\in\mathcal{I}$, where $\mathcal{I}$ is the set of regions. Time is discrete, with annual time steps indexed by $t=0,1,2,\cdots$, where the initial year is the year 2020. All regions are forward-looking with complete information.
@@ 3.1 Emissions
The gross carbon emissions before abatement, in gigatonnes of carbon (GtC), is assumed to be proportional to the gross output for region $i$ at time $t$:
@@
where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}>0$ is the exogenous carbon intensity, calibrated from the projections of GDP and emissions in Ueckerdt et al., (2019). We follow DICE and assume that the amount of emissions abated by each region, $E_{i,t}^{A}$, is expressed as a fraction of the gross emissions, with an emission control rate $\mu_{i,t}\in[0,1]$. [^9] Instead of an emission control rate, researchers also model abatement through reduced use of fossil fuel energy inputs in a production function (see, e.g., Bosetti et al.,, 2006; Bauer et al.,, 2012; Golosov et al.,, 2014; Baldwin et al.,, 2020). However, these models require disaggregation of fossil fuel energy firms, renewable energy firms, final-goods producing firms, and other sectors, so they are often complicated. The emission control rate concept simplifies this. That is,
@@
Thus the emissions net of abatement, $E_{i,t}$, becomes
$E_{i,t}=E_{i}^{\mathrm{Gross}}-E_{i,t}^{A}.$ (11)
We consider a global ETS, or a cap-and-trade system, in which each region is provided with an emission allowance and can trade emission permits with other regions. Regions with emissions exceeding the emission caps can purchase emission permits from other regions. Regions with emissions below their allowance can sell their emission permits to others. The emission cap constraint is represented by
@@
where $\overline{E}_{i,t}$ is the emission cap assigned to region $i$ at time $t$, and $E_{i,t}^{P}$ denotes the amount of emissions traded with other regions. Note that $E_{i,t}^{P}>0$ indicates that region $i$ is a net buyer of emission permits at time $t$, while $E_{i,t}^{P}<0$ implies region $i$ is a net seller of emission permits at time $t$. In a model that does not consider emission permit trade, we simply apply $E_{i,t}^{P}=0$ for all $i$ and $t$. The Paris Agreement establishes upper limits for emissions to ensure that the increase in global temperature remains well below 2°C above pre-industrial levels by the end of this century.
3.2 Climate System
The global average temperature rises as carbon emissions accumulate in the atmosphere. We follow the TCRE climate system and assume that the global average temperature increase above the pre-industrial level is approximately proportional to cumulative global emissions $\mathcal{E}_{t}$, i.e.,
@@ 3.2 Climate System
where the cumulative emissions at 2020, $\mathcal{E}_{0}$, is chosen such that the initial global mean temperature $T_{0}$ is 1.2°C above the pre-industrial level. Here $\zeta=0.0021$ is the contribution rate of CO2 equivalent to temperature, calibrated with the projections of emissions and temperatures in the four Representative Carbon Concentration Pathways (RCPs): RCP 2.6, RCP 4.5, RCP 6.0, and RCP 8.5 (Meinshausen et al.,, 2011). Figure A.1 in Appendix A.3.1 shows that the calibrated TCRE climate system can match all four RCP temperature pathways using their associated RCP emission pathways. Cumulative global emissions evolve according to
@@
3.3 Economic System
A representative consumer in region $i$ has a power utility per period as follows:
$u(c_{i,t})=\frac{c_{i,t}^{1-\gamma}}{1-\gamma},$ (15)
where $c_{i,t}$ is per capita consumption and $\gamma$ is the inverse of intertemporal elasticity of substitution. We follow DICE-2016R Nordhaus, (2017) and set $\gamma=1.45$.
The production technology follows a Cobb-Douglas function of capital and labor inputs. The gross output (in trillion USD), or pre-damage output, $Q_{i,t}$, is given by
$Q_{i,t}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{1-\alpha},$ (16)
where $K_{i,t}$ is capital stock, $L_{i,t}$ is labor, and $\alpha=0.3$ (following Nordhaus, (2017)) is the elasticity of gross output with respect to capital. The total factor productivity $A_{i,t}$ is assumed to follow an exogenous and deterministic time-varying trend, which is calibrated with the GDP projections in Burke et al., (2018). Labor $L_{i,t}$ is also assumed to follow the exogenous population path of region $i$ from the SSP2 scenario (Samir and Lutz,, 2017).
We denote $Y_{i,t}$ as the output net of damages induced by climate changes:
$Y_{i,t}=\frac{1}{1+\pi_{1,i}T_{t}+\pi_{2,i}T_{t}^{2}}Q_{i,t},$ (17)
where $T_{t}$ is the global average temperature increase in °C above the pre-industrial level. We calibrate the climate damage parameters based on the projections of GDP in Burke et al., (2018). We follow DICE and assume that the emission abatement cost is
$\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ (18)
where $b_{1,i,t}=(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))\sigma_{i,t}$. We calibrate the abatement cost parameters, $b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$, by exploiting the projections under different carbon tax scenarios in Ueckerdt et al., (2019), from which we recover the marginal abatement costs and calibrate the parameters we need.
The total consumption of each region under the ETS is obtained by
@@ 3.3 Economic System
where $I_{i,t}$ is investment and $m_{t}$ is the market price of carbon emission permits. The cost of purchasing emission permits from other regions results in reduced consumption, while selling emission permits increases consumption, but this impact is little as the cost or profit from permit trading is small relative to the consumption or output. The capital stock evolves according to the standard neoclassical growth model:
$K_{i,t+1}=(1-\delta)K_{i,t}+I_{i,t},$ (20)
where $\delta=0.1$ is the rate of depreciation of capital stock. We assume that the emission trading market clears in each period. That is,
$\sum_{i\in\mathcal{I}}E_{i,t}^{P}=0,$ (21)
for each time period $t$.
@@ 4.1 The Noncooperative Equilibrium
where $\beta$ is the discount factor. We follow DICE-2016R (Nordhaus,, 2017) and set $\beta=0.985$. Since one region’s emissions will influence the global average temperature, and therefore other regions’ output, the maximization problems of all regions have to be solved simulatenously as a dynamic game. Then we define the dynamic Nash equilibrium of the economy as follows.
@@ 4.2 The Algorithm for the Noncooperative Model
Obtaining the optimal solution of the dynamic model involving multiple regions under noncooperation is challenging. In particular, finding an equilibrium solution for the emission permit prices that satisfies the optimality conditions for each region as well as the market clearing condition poses significant computational challenges. While several studies have incorporated emission prices in their models, no study has successfully solved for the emission permit price in a noncooperative context.
Here we outline the algorithm we develop to obtain the optimal solution for our model. Since $\gamma=1.45$ implies utilities of per capita consumption have negative values and that long-run per capita consumption will be large, the long-run per-capita utility has a small magnitude. Together with the discount factor $\beta=0.985$, the discounted utilities after 300 years are nearly zero and have little impact on the solution in the first 100 years. Let
@@
2. Maximization Step at iteration $j$. For each region $i$, solve the maximization problem (22) without the market clearing condition (21), assuming the permit prices $\{m_{t}^{j-1}:\ t\geq 0\}$ and other regions’ emissions $\{E_{i^{\prime},t}^{j-1}:\ i^{\prime}\neq i,\ t\geq 0\}$ are given from the initialization step when $j=1$ or Step 3 at iteration $j-1$ when $j>1$. The optimal emissions and permits purchased for region $i$ are denoted $\{E_{i,t}^{*,j},E_{i,t}^{P,j}:\ t\geq 0\}$.
@@ 5 Data and Calibration
To calibrate our model with data, we have two objectives. The first objective is to generate regional emission cap pathways ($\overline{E}_{i,t}$) for future periods, constraining the constituent nations of each region to meet their emissions commitments under the Glasgow Pact and the long-term net zero emission targets. The second objective is to determine parameters for total factor productivity ($A_{i,t}$), carbon intensity ($\sigma_{i,t}$), abatement cost ($b_{1,i}$, $b_{2,i}$, $b_{3,i}$, $b_{4,i}$), and climate damage ($\pi_{1,i},\pi_{2,i}$) that reflect the future projections provided in recent studies (Burke et al.,, 2018; Ueckerdt et al.,, 2019; Kahn et al.,, 2021). We obtain these region-specific parameters, which capture regional heterogeneities in GDP growth, emissions, technologies, and climate damages.
Our model’s 12 regions are formed by aggregating 152 countries around the world, following the regional classification in the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ). For the initial year 2020, we obtain country-level historical data on population (billions), capital (\$ trillions, 2020), GDP (\$ trillions, 2020), and emissions (GtC) from the World Bank. For future projections on population growth, we use the SSP2 scenario (Samir and Lutz,, 2017).
@@ 5.1 Regional Emission Cap Pathways
Since there is no global cap-and-trade system or ETS currently in place, we consider the emission commitments outlined in the Paris Agreement and the COP26 Glasgow Climate Pact. Under the Paris Agreement, 195 countries or regions set Nationally Determined Contributions (NDCs) that specify their near-term targets for 2025 or 2030, along with long-term net-zero commitments for 2050 to 2070. These near-term targets were further strengthened during the Glasgow Climate Pact in 2021, with most countries and regions submitting updated or new NDCs. [^10] Individual NDC documents were obtained from the following source: UNFCCC NDC Registry (https://unfccc.int/NDCREG) We collect reports of the most updated Nationally Determined Contributions (NDCs) after the Glasgow Climate Pact and obtain the target years to reach net zero emissions of different countries from Climate Action Tracker. [^11] See https://climateactiontracker.org/. Based on these datasets, we create the baseline regional emission cap pathways for future periods following the strategy detailed below.
@@ 5.3 Carbon Intensity
To obtain the time-varying and region-specific carbon intensities $\sigma_{i,t}$, we use the projections of GDP and emissions in Ueckerdt et al., (2019), who report simulation results of future emissions and GDP under different scenarios based on climate policy regimes, technology portfolios, and carbon tax implementation. As the carbon intensity in our model reflects the zero-carbon tax regime, we employ results from the scenario defined as ‘FFrun451’ in Ueckerdt et al., (2019). Specifically, this FFrun451 scenario corresponds to weak policy baseline before Paris Agreement with limited renewable technology portfolio and zero carbon tax. Based on the equation (9), we calculate the carbon intensities as $\sigma_{i,t}=E_{i,t,\mathrm{FFrun451}}^{U}/Q_{i,t,\mathrm{FFrun451}}$, where $E_{i,t,\mathrm{FFrun451}}^{U}$ and $Q_{i,t,\mathrm{FFrun451}}$ are the projected regional emissions and GDP under the FFrun451 scenario for region $i$ at time $t$.  [^15] Ueckerdt et al., (2019) provide data for 11 regions. Upon comparison, we find that the countries constituting the ‘Rest of the World (ROW)’ region are the ones that are in the ‘Other High Income (OHI)’ and ‘Eurasia’ region in our study. Therefore, the carbon intensity obtained from the ROW in Ueckerdt et al., (2019) corresponds to that of OHI and Eurasia regions in our work.

### v4

@@ 4 A Dynamic Regional Model of Climate and the Economy with an ETS
We now introduce a dynamic regional model of climate and economy that integrates a global ETS across multiple regions. Our model framework extends the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ) by incorporating the global ETS in a dynamic setting and the TCRE climate system with annual time steps. One of the main focuses of the model is to characterize the equilibrium path of carbon prices in a noncooperative environment, where a social planner of each region maximizes its own social welfare by optimally choosing emissions abatement, permit purchases in the global carbon market, and consumption over time. Future regional emissions are constrained by emission caps based on commitments established under the Paris Agreement, later updated by the Glasgow Pact, and by net-zero targets.
The macroeconomic framework of our model employs a multi-regional representation of the Ramsey growth model. Each region is indexed by $i\in\mathcal{I}$, where $\mathcal{I}$ is the set of regions. Time is discrete and infinite, with annual time steps indexed by $t=0,1,2,\cdots$. All agents are forward-looking with complete information. The model presented here abstracts from uncertainty, excludes international trade of goods other than emission permits, and assumes frictionless trading in the ETS.
4.1 The Economic System
Each region consists of a representative household with a population size of $L_{i,t}$, and utility of the representative consumer is given by
$u(c_{i,t})=\frac{c_{i,t}^{1-\gamma}}{1-\gamma},$ (3)
where $c_{i,t}$ is per capita consumption and $\gamma$ is the inverse of intertemporal elasticity of substitution. Following DICE-2016R (Nordhaus,, 2017), $\gamma$ is set to 1.45.
There is a representative firm in each region that employs a Cobb-Douglas production technology using capital and labor as inputs, and produces a numeraire good whose price is normalized to 1. The representative household owns all input factors and the firm of the regional economy. The gross output, or pre-damage output, $Q_{i,t}$, is given by
$Q_{i,t}=A_{i,t}K_{i,t}^{\alpha}L_{i,t}^{1-\alpha},$ (4)
where $K_{i,t}$ is capital stock, and $\alpha=0.3$ is the elasticity of gross output with respect to capital, as in Nordhaus, (2017). Consistent with the standard neoclassical growth model, capital depreciates over time, and the firm invests to replenish and accumulate capital stock. The evolution of capital follows
$K_{i,t+1}=(1-\delta)K_{i,t}+I_{i,t},$ (5)
where $I_{i,t}$ is investment and $\delta=0.1$ is the rate of depreciation of capital stock.
Regions experience climate damages resulting from the externalities of global emissions, with the extent of these damages varying across regions. The output net of climate-induced damages, denoted as $Y_{i,t}$, is given by:
$Y_{i,t}=\left(\frac{1}{1+\pi_{1,i}T_{t}+\pi_{2,i}T_{t}^{2}}\right)Q_{i,t},$ (6)
where $\pi_{1,i}$ and $\pi_{2,i}$ are region-specific climate damage parameters, and $T_{t}$ is the global average temperature increase in degrees Celsius above the pre-industrial level. The specification captures the adverse (or potentially beneficial for some regions) effects of rising global average temperature, where local damages increase nonlinearly with temperature, following the quadratic form commonly used in the climate economics literature (Nordhaus,, 2014, 2017).
@@ 4.2 Emissions
Each region’s economic activity produces carbon emissions, and the representative firm faces emission constraints of the region. The gross emissions before abatement, in gigatonnes of carbon (GtC), is assumed to be proportional to the gross output for region $i$ at time $t$:
@@
where $Q_{i,t}$ is the gross output, and $\sigma_{i,t}>0$ is the exogenous carbon intensity. The firm may choose to reduce a fraction of gross emissions, with its efforts represented by emission control rate $\mu_{i,t}\in[0,1]$. [^8] Alternatively, abatement can also be modeled through the reduced use of fossil fuel energy inputs in the production function (see, e.g., Bosetti et al.,, 2006; Bauer et al.,, 2012; Golosov et al.,, 2014; Baldwin et al.,, 2020). However, such models often require disaggregation of fossil fuel energy firms, renewable energy firms, final-goods producers, and other sectors, increasing the complexity of the model. The emission control rate approach simplifies this by offering a more streamlined representation. The amount of emissions abated by each region, $E_{i,t}^{A}$, is then expressed as follows.
@@
The emission abatement efforts incur costs to the firm and are heterogeneous across regions due to technological differences. The emission abatement cost, $\Phi_{i,t}$, is specified as
$\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ (9)
where $b_{1,i,t}=(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))\sigma_{i,t}$. The parameters $b_{1,i}$, $b_{2,i}$, $b_{3,i},$ and $b_{4,i}$ govern the cost structure of emission abatement cost, $\Phi_{i,t}$, which depends on both the gross output $Q_{i,t}$ and the exponential function of the emission control rate, $\mu_{i,t}$. This specification captures the dynamic nature of abatement costs, reflecting potential technological advancements that reduce the cost of emissions abatement over time.
In a global ETS, or a cap-and-trade system, each region is provided with an emission allowance and can trade emission permits with other regions. The representative firm in each region that emits beyond its cap can purchase permits, while those emitting below their allowance can sell excess permits. Denoting the emissions net of abatement (net emissions) as $E_{i,t}=E_{i}^{\mathrm{Gross}}-E_{i,t}^{A}$, the emission cap constraint is represented by
@@
where $\overline{E}_{i,t}$ is the emission cap assigned to region $i$ at time $t$, and $E_{i,t}^{P}$ denotes the amount of emissions purchased from other regions. Note that $E_{i,t}^{P}>0$ indicates that region $i$ is a net buyer of emission permits at time $t$, while $E_{i,t}^{P}<0$ implies region $i$ is a net seller of emission permits at time $t$. In a model that does not consider emission permit trade, the emissions purchase is simply set at $E_{i,t}^{P}=0$ for all $i$ and $t$.
4.3 The Climate System
The global average temperature rises as carbon emissions accumulate in the atmosphere. Adopting the TCRE climate system representation (Matthews et al.,, 2009), it is assumed that the global average temperature increase above the pre-industrial level is approximately linear to cumulative global emissions $\mathcal{E}_{t}$, i.e.,
@@ 4.3 The Climate System
where $\zeta$ represents the contribution rate of cumulative global emissions to temperature. The cumulative global emissions evolve according to
@@
This dynamic process captures the accumulation of emissions in the atmosphere over time, with each region contributing to the global emissions stock and, consequently to the increase in global temperature.
4.4 Market Clearing
The goods market clearing implies that the total consumption of each region under the ETS is constrained by
@@ 4.4 Market Clearing
where $m_{t}$ is the market equilibrium price of emission permits. As in the static framework, the last term on the right-hand side reflects the cost or revenue generated from the emission permit trade.
Finally, the emission trading market clears each period, given by: [^9] The current model assumes that intertemporal lending or borrowing of emission permits is not allowed.
$\sum_{i\in\mathcal{I}}E_{i,t}^{P}=0.$ (14)
If it is a partial ETS and region $i$ does not participate in the partial ETS, then we set $E_{i,t}^{P}$ to be fixed at zero.
@@ 5.1 The Noncooperative Equilibrium
where $\beta$ is the discount factor. We follow DICE-2016R (Nordhaus,, 2017) and set $\beta=0.985$. Since one region’s emissions will influence the global average temperature, and therefore other regions’ output, the maximization problems of all regions have to be solved simultaneously as a dynamic game. Then we define the dynamic Nash equilibrium of the economy as follows.
@@ 5.2 The Algorithm for the Noncooperative Model
Obtaining the optimal solution of the dynamic model involving multiple regions under noncooperation is challenging. In particular, finding an equilibrium solution for the emission permit prices that satisfies the optimality conditions for each region as well as the market clearing condition poses significant computational challenges.
Here we outline the algorithm we develop to obtain the optimal solution for our model. With the discount factor $\beta=0.985$, the discounted utilities after 300 years are nearly zero and have little impact on the solution in the first 100 years. Let
@@
2. Maximization Step at iteration $j$. For each region $i$, solve the maximization problem (15) without the market clearing condition (14), assuming the permit prices $\{m_{t}^{j-1}:\ t\geq 0\}$ and other regions’ emissions $\{E_{i^{\prime},t}^{j-1}:\ i^{\prime}\neq i,\ t\geq 0\}$ are given from the initialization step when $j=1$ or Step 3 at iteration j1j-1 when $j>1$. The optimal emissions and permits purchased for region $i$ are denoted $\{E_{i,t}^{*,j},E_{i,t}^{P,j}:\ t\geq 0\}$.
@@ 6 Data and Calibration
When taking our model to data, two key objectives are pursued. The first objective is to generate regional emission cap pathways ($\overline{E}_{i,t}$) for future periods, constraining the constituent nations of each region to meet their emissions commitments under the Glasgow Pact and the long-term net zero emission targets. The second objective is to determine parameters for total factor productivity ($A_{i,t}$), carbon intensity ($\sigma_{i,t}$), abatement cost ($b_{1,i}$, $b_{2,i}$, $b_{3,i}$, $b_{4,i}$), and climate damage ($\pi_{1,i},\pi_{2,i}$) that reflect the future projections provided in recent studies (Burke et al.,, 2018; Ueckerdt et al.,, 2019; Kahn et al.,, 2021). We obtain these region-specific parameters, which capture regional heterogeneities in GDP growth, emissions, technologies, and climate damages.
We assume that the world is divided into 12 aggregated regions: the United States (US), the EU, Japan, Russia, Eurasia, China, India, Middle East (MidEast), Africa, Latin America (LatAm), Other High-Income countries (OHI) and other non-OECD Asia (OthAs). These 12 regions are formed by aggregating 152 countries around the world, following the regional classification in the RICE model (Nordhaus and Yang,, 1996; Nordhaus, 2010b, ). [^10] See Appendix 1 A.2 for the full list of countries and regional aggregation. Note that, while we present our simulation results at a regionally aggregated level for computational tractability, our methods are applicable to a larger number of regions and can be readily extended to obtain country-level outcomes. For example, the regional SCC for Africa can be disaggregated to approximate country-level SCCs, since the regional SCC is equal to the sum of the SCCs of the individual countries within the region.  [^11] Regional SCC is the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period. The initial year is set to 2020, with country-level historical data on population (billions), capital (\$ trillions, 2020), GDP (\$ trillions, 2020), and emissions (CO2 equivalent, GtC) sourced from the World Bank.  [^12] To obtain initial capital stock, we used capital formation data from the World Bank. Specifically, we computed $K_{i,2020}=\sum_{g=0}^{20}(0.9)^{g}\text{CF}_{i.2020-g}$, where $\text{CF}_{i,t}$denotes capital formation adjusted to 2020 values in trillions of USD. For future population pathways, projections from the SSP2 scenario (Samir and Lutz,, 2017) are employed.
@@ 6.1 Regional Emission Cap Pathways
Since there is no global cap-and-trade system or ETS currently in place, we consider the emission commitments outlined in the Paris Agreement and the COP26 Glasgow Climate Pact. Under the Paris Agreement, 195 countries or regions set Nationally Determined Contributions (NDCs) that specify their near-term targets for 2025 or 2030, along with long-term net-zero commitments for 2050 to 2070. These near-term targets were further strengthened during the Glasgow Climate Pact in 2021, with most countries and regions submitting updated or new NDCs. [^13] Individual NDC documents were obtained from the following source: UNFCCC NDC Registry (https://unfccc.int/NDCREG) We collect reports of the most updated NDCs after the Glasgow Climate Pact and obtain the target years to reach net zero emissions of different countries from Climate Action Tracker. [^14] See https://climateactiontracker.org/. Based on these datasets, we create the baseline regional emission cap pathways for future periods following the strategy detailed below.
@@ 6.3 Carbon Intensity
To obtain the time-varying and region-specific carbon intensities $\sigma_{i,t}$, we use the projections of GDP and emissions in Ueckerdt et al., (2019), who report simulation results of future emissions and GDP under different scenarios based on climate policy regimes, technology portfolios, and carbon tax implementation. As the carbon intensity in our model reflects the zero-carbon tax regime, we employ results from the scenario defined as ‘FFrun111’ in Ueckerdt et al., (2019). Specifically, this FFrun111 scenario corresponds to climate action from 2010 with full technology portfolio and no carbon tax. . Based on the equation (7), we calculate the carbon intensities as $\sigma_{i,t}=E_{i,t,\mathrm{FFrun111}}^{U}/Q_{i,t,\mathrm{FFrun111}}$, where $E_{i,t,\mathrm{FFrun111}}^{U}$ and $Q_{i,t,\mathrm{FFrun111}}$ are the projected regional emissions and GDP under the FFrun111 scenario for region $i$ at time $t$.  [^18] Ueckerdt et al., (2019) provide data for 11 regions. Upon comparison, we find that the countries constituting the ‘Rest of the World (ROW)’ region are the ones that are in the ‘Other High Income (OHI)’ and ‘Eurasia’ region in our study. Therefore, the carbon intensity obtained from the ROW in Ueckerdt et al., (2019) corresponds to that of OHI and Eurasia regions in our work.